# A new paradigm of mortality modeling via individual vitality dynamics

### v2

@@ 4.5 Cause-of-death modeling
$\mathbb{E}_{v}\left[\mathrm{e}^{-q\tau_{J}}\mathbf{1}\left(\tau_{J}<\tau_{Y}\right)\right]=$ $\int_{0}^{t_{v}^{*}}\lambda\mathrm{e}^{-(q+\lambda)t-\alpha b^{*}_{v}}\cdot I_{0}\left(2\sqrt{\alpha\lambda tb^{*}_{v}}\right)\mathrm{d}t,$ (13)
@@
$\mathbb{E}_{v}\left[\mathrm{e}^{-q\tau_{Y}}\mathbf{1}\left(\tau_{Y}<\tau_{J}\right)\right]=$ $\int_{0}^{t^{*}_{v}}\mathrm{e}^{-\left(q+\lambda\right)t-\alpha b^{*}_{v}}\cdot bc^{x+t}\sqrt{\frac{\alpha\lambda t}{b^{*}_{v}}}I_{1}\left(2\sqrt{\alpha\lambda tb^{*}_{v}}\right)\mathrm{d}t+\mathrm{e}^{-(q+\lambda)t^{*}_{v}},$ (14)
where $t^{*}_{v}=\frac{\ln(bc^{x}+v\ln c)-\ln b}{\ln c}-x$ (which refers to the remaining lifetime without any accidents/jumps), $b^{*}_{v}=v-bc^{x}\left(c^{t}-1\right)/\ln c$, and
@@
$\mathrm{Pr}_{v}\left(\tau_{J}\in\mathrm{d}t,\tau_{J}<\tau_{Y}\right)$ $=\mathbb{1}(t<t^{*})\lambda\mathrm{e}^{-\alpha b^{*}_{v}-\lambda t}I_{0}\left(2\sqrt{\lambda\alpha tb^{*}_{v}}\right)\mathrm{d}t,$
@@
$\mathrm{Pr}_{v}\left(\tau_{Y}\in\mathrm{d}t,\tau_{Y}<\tau_{J}\right)=$ $\mathbb{1}(t<t^{*})\mathrm{e}^{-\lambda t-\alpha b^{*}_{v}}\cdot bc^{x+t}\sqrt{\frac{\alpha\lambda t}{b^{*}_{v}}}I_{1}\left(2\sqrt{\alpha\lambda tb^{*}_{v}}\right)\mathrm{d}t+\mathrm{e}^{-\lambda t}\delta_{t^{*}}(\mathrm{d}t),$
@@ 5 A Numerical Illustration
In this section, we provide an illustrative example to discuss the numerical estimation procedure of the proposed vitality-based mortality modeling framework. We consider the jump-diffusion specification shown in Section 3.1.3. Recall from equations (4) and (5) that the dynamics of vitality is governed by
$V(t)=V(0)-\int_{0}^{t}\mu_{x}(s)\mathop{}\!\mathrm{d}s-\sigma B(t)-\sum_{i=1}^{N(t)}Z_{i},$
and the resulting $T$-year survival probability is
$\Pr(\tau>T)=\int_{0}^{\infty}\mathbb{P}\left(v-\sigma B(t)-\sum_{i=1}^{N(t)}Z_{i}\geq\int_{0}^{t}\mu_{x}(s)\mathop{}\!\mathrm{d}s,\;\forall t\leq T\right)\mathop{}\!\mathrm{d}F_{0}(v).$
5.1 Estimation simplifications
The computational requirements for evaluating the conditional probability in the equation of $\Pr(\tau>T)$ are quite intensive. For ease of computation, we make the following assumptions to achieve an approximation:
@@ 5.1 Estimation simplifications
$Y(t)=a_{i}(t-\lfloor t\rfloor)+b_{i},$
for $i\leq t<i+1$ and $i\in\{0,1,2,\ldots\}$, where
$b_{i}$ $=\int_{0}^{i}\mu_{x}(s)\mathop{}\!\mathrm{d}s\quad\text{and}\quad a_{i}=\int_{i}^{i+1}\mu_{x}(s)\mathop{}\!\mathrm{d}s.$
Using the well-known formula for the non-crossing probability for a linear boundary $at+b$ before time $s$, that is
$\Pr(B(t)<at+b\text{ for all }t\leq s\big{|}B(s)=x)=1-\exp\left\{-\frac{2b(as+b-x)}{s}\right\},$
together with the fatal jump assumption, the conditional survival probability is given by
${}_{k}p_{x}(v)=\mathbb{E}\left[\left(\prod_{s=1}^{k}\mathbb{1}_{B(s)<\frac{v-Y(s)}{\sigma}}\cdot\left(1-e^{-\frac{2(v-Y(s-1)-\sigma B(s-1))(v-Y(s)-\sigma B(s))}{\sigma^{2}}}\right)\right)\cdot\exp\left(-\int_{0}^{k}\lambda(s)\mathop{}\!\mathrm{d}s\right)\right],k\in\mathbb{N}^{+},$
for a given initial vitality $v>0$. It follows that the $T$-year survival probability becomes
@@
5.2 Maximum likelihood estimation
In this subsection, we consider the maximum likelihood estimation method for estimating the parameters of the vitality model. We begin by describing the likelihood function for a cohort mortality dataset.
Consider a cohort of individuals all aged $x$ at time $t=0$. Let $E_{x}$ be the initial exposure at time 00, and let $D_{x+t}$ be the observed death counts at age $x+t$ for $t\in\{0,1,2,\ldots\}$. We assume that $D_{x+t}$ follows a multinomial distribution, and thus the log-likelihood function of the cohort mortality data is
@@ 5.2 Maximum likelihood estimation
Given the complexity of our vitality model, relying solely on a cohort mortality dataset from a national population is inadequate for estimation purposes. Specifically, a national cohort dataset does not distinguish between deaths caused by natural aging and those resulting from accidents. This limitation makes it challenging to differentiate between accidental deaths, captured by the jump component, and natural deaths, represented by the initial and depletion components. In terms of parameter estimation, the intensity rate $\lambda(t)$ of the Poisson process $N(t)$ cannot be easily estimated without information that explicitly categorizes accidental deaths.
To address this issue, we utilize mortality data specifically related to external causes of death. In particular, we use external death counts from the Centers for Disease Control and Prevention (CDC) of the United States to estimate the intensity rate $\lambda(t)$. The remaining parameters, $\theta_{-\lambda}=\theta\backslash\{\lambda(t)\}_{t=0,1,\cdots}$, are then estimated using the maximum likelihood estimation method. This approach allows us to accurately separate the jump component from the rest of the components. The numerical results of our estimation procedures are presented in the next subsection.
5.3 Numerical estimation results
To estimate the vitality model, we utilize the U.S. male mortality data from the Human Mortality Database (HMD) [^1] Human Mortality Database. Max Planck Institute for Demographic Research (Germany), University of California, Berkeley (USA), and French Institute for Demographic Studies (France). Available at www.mortality.org.. We focus on the cohort born in 1910, with ages 60 and above during the period from 1970 to 2016. Additionally, we employ data from the Centers for Disease Control and Prevention (CDC) that consists of the average accidental death rates for age groups in ten-year intervals, spanning from 1968 to 2016. The jump intensity $\lambda(t)$ is directly determined as a constant within each age group based on the CDC data. The remaining parameters are fitted to the HMD data described above.
Recall that the jump-diffusion vitality model from Section 3.1.3 uses the Gompertz law to specify the trend component; that is, $\mu_{x}(t)=bc^{x+t}$. It follows that the vitality dynamics are governed by
@@ 5.3 Numerical estimation results
where $V(0)\sim\text{Exp}(1)$. Given that the intensity rate of $N(t)$ has been externally estimated and we assumed $Z_{i}=\infty$ for all $i$, the remaining parameters to be estimated are $b$, $c$, and $\sigma$. In addition to the vitality model, we also estimate the parameters of the Gompertz law for comparison purposes.
Table 1 reports the estimated values of $b$, $c$, and $\sigma$ for the vitality model and the estimated values of $b$ and $c$ for the Gompertz law. The maximized log-likelihood of each model is also reported in the last row. We observe that the estimated values of $b$ and $c$ are similar between the vitality model and the Gompertz law, indicating that both models accurately describe the underlying mortality curve via the assumed functional form, $\mu_{x}(t)=bc^{x+t}$. In addition to the Gompertz law, the vitality model has an estimated value of 0.0036 for $\sigma$, which reflects the random fluctuations in vitality over the lifetime of the cohort. Overall, the vitality model achieves a higher maximized log-likelihood of -2,181 compared to the Gompertz law’s -2,591.
@@
$b$ 0.00015391 0.0001744
$c$ 1.0834 1.082
$\sigma$ 0.0036
$l(\theta)$2,181 -2,591
Table 1: Parameter estimates and maximized log-likelihood for the vitality model and the Gompertz law.
Figure 2 shows the fitted values of death rates and survival probabilities for ages 60-100 and periods 0-40 years under the vitality model and the Gompertz law. The observed values from the empirical data are also plotted for comparison. We observe that both the vitality model and the Gompertz law are able to depict a well-fitted mortality curve for the cohort considered and also produce reasonable survival probabilities at age 60 for the next 40 years. Comparing the two models, the vitality model appears to better fit the cohort’s mortality data and produce a mortality curve that is closer to the observed values.
Figure 2: Fitted log death rates $\ln q_{x}$ for $x\in(60,100)$ and survival probabilities ${}_{t}p_{60}$ for $t\in(0,40)$ under the vitality model (blue solid line) and the Gompertz law (red dash line). The observed values are depicted by dots.
@@ 6 Conclusion
In this paper, we introduced a vitality-based modeling framework that offers a new perspective on mortality modeling. The proposed framework integrates multiple components to capture the complexities of vitality dynamics over an individual’s lifetime. By incorporating the initial vitality level, natural depletion trends, stochastic diffusion processes, and sudden jumps, our framework provides a comprehensive and flexible approach to mortality modeling. The adaptability of the four components allows for the incorporation of various factors affecting mortality, providing deeper insights into population mortality trends and individual health trajectories.
@@
Through both static and dynamic specifications of the proposed framework, we demonstrated its versatility in addressing various mortality modeling scenarios. Our framework not only replicates traditional mortality laws but also extends to more advanced stochastic mortality models, offering a unified structure for understanding mortality dynamics. An illustrative example is provided to showcase how a vitality model can be fitted using the maximum likelihood estimation method. Additionally, the framework’s modularity ensures that it can be tailored to specific research needs, making it a valuable tool for both theoretical exploration and practical applications.
@@
Our findings suggest that the vitality-based modeling framework offers significant practical applications, such as life insurance valuation, disability modeling, cause-of-death modeling, explaining subjective beliefs about mortality, and deriving optimal consumption solutions. By capturing the initial level, natural aging, inherent randomness, and sudden changes in an individual’s vitality, the framework supports a better interpretation of mortality dynamics in the aforementioned applications. Lastly, we emphasize that the proposed framework can be further utilized in more nuanced scenarios to improve results and decision-making in various fields.
@@
While we used four components to develop a comprehensive modeling framework for vitality dynamics, the specific structures considered for each component are not exhaustive in this paper. There are numerous possible extensions for articulating the four components. For instance, one could consider a multivariate stochastic process to govern the diffusion component, reflecting the fact that multiple underlying forces may drive the randomness in an individual’s vitality. The initial component could be modified to measure and explain healthy life expectancy. The jump component can also be expanded to consider jumps categorized by different causes of death, such as natural disasters, car accidents, heart attacks, and suicide.
@@
We have demonstrated in this paper that the proposed modeling framework is capable of replicating mortality laws and stochastic mortality models. Beyond these existing models, further exploration can be conducted to show that continuous-time intensity-based models, such as the Vasicek process applied to the force of mortality, can also be mimicked by a vitality model. It would be interesting to understand how the first passage time problem connects with an continuous-time mortality model to produce explicit solutions for survival probabilities. This connection could potentially provide new insights and more accurate survival probability estimates.
@@
Although we have explored various applications of the vitality-based modeling framework, there are several other potential applications that we did not focus on. For example, a dynamic vitality model can be used to manage longevity risk by developing dynamic hedging strategies and pricing longevity-linked securities. Additionally, since our framework models individual vitality, it can be employed to measure the basis risk between an individual insurance product and an index-based security. These extensions would provide further advancements to capital market solutions for longevity risk management.
@@
Lastly, we acknowledge that the vitality dynamics of an individual cannot be directly observed, or at least not easily. To accurately estimate a vitality model using data beyond the national population data presented in this paper, one would need to use personal data over time or clinical data on vital measures. Potential sources of such data could include personal medical devices, such as smartwatches or mobile health trackers, which continuously monitor health metrics like heart rate, physical activity, and sleep patterns. Integrating these diverse data sources would allow for a more precise estimation of vitality dynamics. We remark that this research direction is highly related to InsurTech and could inspire future studies on personalized insurance products.
@@ Acknowledgement
We extend our gratitude for the helpful and constructive comments received at the School of Actuarial Science at Université Laval, the 27th International Congress on Insurance: Mathematics and Economics, and the 59th Actuarial Research Conference. Xiaobai Zhu acknowledges the support from the Research Grants Council of the Hong Kong Special Administrative Region, China (CUHK 24615523), and National Natural Science Foundation of China (No. 12301613). Zijia Wang acknowledges the support from a start-up grant from The Chinese University of Hong Kong.
Appendix A Proofs
A.1 Death Plateau Models
@@
The remaining steps follow exactly the same as before, and we will end up with the same survival function as in Appendix A.1.
@@
If $a=1$, $p=\alpha$, and $b=\frac{v}{\phi}$, then we recover the same survival function as in Appendix A.1.
@@ A.2 Alternative Expression of Mortality Laws
A.3 Dynamic Model

### v3

@@ 4.5 Cause-of-death modeling
$\mathbb{E}_{v}\left[\mathrm{e}^{-q\tau_{J}}\mathbb{1}\left(\tau_{J}<\tau_{Y}\right)\right]=$ $\int_{0}^{t_{v}^{*}}\lambda\mathrm{e}^{-(q+\lambda)t-\alpha b^{*}_{v}(t)}\cdot I_{0}\left(2\sqrt{\alpha\lambda tb^{*}_{v}(t)}\right)\mathrm{d}t,$ (13)
@@
$\mathbb{E}_{v}\left[\mathrm{e}^{-q\tau_{Y}}\mathbb{1}\left(\tau_{Y}<\tau_{J}\right)\right]=$ $\int_{0}^{t^{*}_{v}}\mathrm{e}^{-\left(q+\lambda\right)t-\alpha b^{*}_{v}(t)}\cdot bc^{x+t}\sqrt{\frac{\alpha\lambda t}{b^{*}_{v}(t)}}I_{1}\left(2\sqrt{\alpha\lambda tb^{*}_{v}(t)}\right)\mathrm{d}t+\mathrm{e}^{-(q+\lambda)t^{*}_{v}},$ (14)
where $t^{*}_{v}=\frac{\ln(bc^{x}+v\ln c)-\ln b}{\ln c}-x$ (which refers to the remaining lifetime without any accidents/jumps), $b^{*}_{v}(t)=v-bc^{x}\left(c^{t}-1\right)/\ln c$, and
@@
$\mathrm{Pr}_{v}\left(\tau_{J}\in\mathrm{d}t,\tau_{J}<\tau_{Y}\right)$ $=\mathbb{1}(t<t^{*})\lambda\mathrm{e}^{-\alpha b^{*}_{v}(t)-\lambda t}I_{0}\left(2\sqrt{\lambda\alpha tb^{*}_{v}(t)}\right)\mathrm{d}t,$
@@
$\mathrm{Pr}_{v}\left(\tau_{Y}\in\mathrm{d}t,\tau_{Y}<\tau_{J}\right)=$ $\mathbb{1}(t<t^{*})\mathrm{e}^{-\lambda t-\alpha b^{*}_{v}(t)}\cdot bc^{x+t}\sqrt{\frac{\alpha\lambda t}{b^{*}_{v}(t)}}I_{1}\left(2\sqrt{\alpha\lambda tb^{*}_{v}(t)}\right)\mathrm{d}t+\mathrm{e}^{-\lambda t}\delta_{t^{*}}(\mathrm{d}t),$
@@ 5 A Numerical Illustration
In this section, we provide an illustrative example on the numerical estimation procedure of the proposed framework. We consider the jump-diffusion vitality model discussed in Section 3.1.3. Recall that the vitality process $V(t)$ is described by equation (4) while the resulting survival probability $\Pr(\tau>T)$ is given in equation (5).
5.1 Simplifications
The computational requirements for evaluating equation (5) are quite intensive. For ease of computation, we make the following assumptions to achieve an approximation:
@@ 5.1 Simplifications
$Y(t)=\alpha_{i}(t-\lfloor t\rfloor)+\beta_{i},$
for $i\leq t<i+1$ and $i\in\{0,1,2,\ldots\}$, where $\beta_{i}=\int_{0}^{i}\mu_{x}(s)\mathop{}\!\mathrm{d}s$ and $\alpha_{i}=\int_{i}^{i+1}\mu_{x}(s)\mathop{}\!\mathrm{d}s$. Using the well-known formula for the non-crossing probability of a linear boundary $\alpha t+\beta$ before time $s$; that is,
$\Pr\left(B(t)<\alpha t+\beta\text{ for all }t\leq s\big{|}B(s)=x\right)=1-\exp\left\{-\frac{2\beta(\alpha s+\beta-x)}{s}\right\},$
together with the fatal jump assumption, the survival probability conditional on an initial vitality $v>0$ is given by
${}_{k}p_{x}(v)=\mathbb{E}\left[\left(\prod_{s=1}^{k}\mathbb{1}_{B(s)<\frac{v-Y(s)}{\sigma}}\left(1-e^{-\frac{2(v-Y(s-1)-\sigma B(s-1))(v-Y(s)-\sigma B(s))}{\sigma^{2}}}\right)\right)\exp\left(-\int_{0}^{k}\lambda(s)\mathop{}\!\mathrm{d}s\right)\right],$
for $k\in\mathbb{N}^{+}$. It follows that the $T$-year survival probability becomes
@@
5.2 Model estimation
Given the complexity of vitality mortality models, using solely the mortality data from a national population is inadequate for estimation purposes. Specifically, such dataset does not distinguish between deaths caused by natural aging and those resulting from accidents. This limitation makes it challenging to differentiate between accidental deaths, captured by the jump component, and natural deaths, represented by the initial and trend components. In terms of parameter estimation, the intensity rate $\lambda(t)$ of the Poisson process $N(t)$ cannot be easily estimated without information that explicitly categorizes accidental deaths.
To address this issue, we utilize mortality data that are specifically related to accidental causes of death. In particular, we use the WONDER dataset from the Centers for Disease Control and Prevention (CDC) to determine the intensity rate $\lambda(t)$. The remaining parameters, $\theta_{-\lambda}=\theta\backslash\{\lambda(t)\}_{t=0,1,\cdots}$, are then estimated using the maximum likelihood estimation method. This approach allows us to separate the jump component from the rest of the components. More specifically, we estimate $\theta_{-\lambda}$ using the U.S. male mortality data from the Human Mortality Database (HMD) [^1] Human Mortality Database. Max Planck Institute for Demographic Research (Germany), University of California, Berkeley (USA), and French Institute for Demographic Studies (France). Available at www.mortality.org.. We focus on the cohort born in 1910, with age 60 and above during the period of 1970 to 2016. We estimate $\lambda(t)$ using the average accidental death rates for age groups in ten-year intervals during the period of 1968 to 2016 from the CDC WONDER dataset.
The likelihood function of our maximum likelihood estimation is given as follows. Consider a cohort of individuals all aged $x$ at time $t=0$. Let $E_{x}$ be the initial exposure at time 00 and $D_{x+t}$ be the observed death counts at age $x+t$ for $t\in\mathbb{N}^{+}$. We assume that $D_{x+t}$ follows a multinomial distribution, which leads to the log-likelihood function as
@@ 5.2 Model estimation
5.3 Numerical results
Using the Gompertz law to specify the trend component of the vitality dynamics, the jump-diffusion vitality model is specified as
@@ 5.3 Numerical results
where $V(0)\sim\text{Exp}(1)$. Given that the intensity rate of $N(t)$ has been externally determined and we assumed $Z_{i}=\infty$ for all $i$, the remaining parameters to be estimated are $b$, $c$, and $\sigma$. In addition to the vitality model, we also estimate the parameters of the Gompertz law for comparison purposes.
Table 1 reports the estimated values of $b$, $c$, and $\sigma$ from the vitality model and the Gompertz law for both males (left panel) and females (right panel). The maximized log-likelihood of each model is reported at the last row of the corresponding table. We observe that the estimated values of $b$ and $c$ are similar between the vitality model and the Gompertz law, indicating that both models accurately describe the underlying mortality curve via their assumed functional form (i.e., $\mu_{x}(t)=bc^{x+t}$). Compared to the Gompertz law, the vitality model additionally has an estimated value of $\sigma$ at 0.0098 and 0.0034 for males and females, respectively. These values reflect the random fluctuations in vitality over the lifetime of the corresponding cohort of males and females. Overall, the vitality model achieved a higher maximized log-likelihood than the Gompertz law for both genders.
@@
$b$ $1.4819\times 10^{-4}$ $1.7440\times 10^{-4}$
$c$ 1.0838 1.0820
$\sigma$ $9.8267\times 10^{-3}$
$l(\theta)$2,057 -2,591
Model Vitality Gompertz
$b$ $1.8107\times 10^{-5}$ $1.9915\times 10^{-5}$
$c$ 1.1058 1.1048
$\sigma$ $3.4139\times 10^{-3}$
$l(\theta)$3,375 -3,901
Table 1: Parameter estimates and maximized log-likelihood of the vitality model and the Gompertz law for males (left) and females (right).
@@ 6 Conclusion
In this paper, we introduced a vitality-based modeling framework that offers a new perspective on mortality modeling. The proposed framework integrates four unique components to capture the complexities of vitality dynamics over an individual’s life course. The adaptability of the four components allows for the incorporation of various factors affecting mortality, providing deeper insights into population mortality trends and individual health trajectories.
@@
Through both static and dynamic parameter specifications of the proposed framework, we demonstrated its versatility in addressing various mortality modeling scenarios. Our framework not only replicates traditional mortality laws but also extends to more advanced stochastic mortality models, offering a unified structure for understanding mortality dynamics. An illustrative example is provided to showcase how a vitality model can be fitted using the maximum likelihood estimation method.
@@
Our investigation suggests that the vitality-based modeling framework offers significant practical applications, such as life insurance valuation, disability modeling, cause-of-death modeling, explaining subjective beliefs on mortality, and deriving optimal consumption solutions. We emphasize that the proposed framework can be further utilized in more nuanced scenarios to improve results and decision-making in various fields.
@@
While we used four components in the proposed framework, the specific structures considered for each component are not exhaustive in this paper. There are numerous possible extensions for articulating the four components. For instance, one could consider a multivariate stochastic process to govern the diffusion component, reflecting the fact that multiple underlying forces may drive the randomness in an individual’s vitality. The jump component can also be expanded to consider jumps categorized by different causes of death.
@@
We have demonstrated in this paper that the proposed framework is capable of replicating mortality laws and factor-based models. Further exploration can be conducted to show that intensity-based models, such as the Vasicek process applied to the force of mortality, can also be mimicked by the proposed framework. It would be interesting to understand how the first passage time problem connects with an continuous-time intensity-based model. This connection could potentially provide new insights and more accurate survival probability estimates.
@@
There are several other potential applications that we did not consider in this paper. For instance, a vitality model with dynamic parameters can be used to manage longevity risk by developing hedging strategies and pricing longevity-linked securities. Additionally, since our framework models individual vitality, it can be employed to measure the basis risk between an individual insurance product and an index-based security. These extensions would provide further advancements to capital market solutions for longevity risk management.
@@
Lastly, we acknowledge that the vitality dynamics of an individual cannot be directly observed and easily estimated. To improve the estimation of a vitality model, one would need to use personal health data or clinical data on vital measures. Potential sources of such data could include personal medical devices, such as smartwatches or mobile health trackers. We remark that this research direction is highly related to InsurTech and could inspire future studies on personalized insurance products.
@@ Acknowledgement
Xiaobai Zhu acknowledges the support from the Research Grants Council of the Hong Kong Special Administrative Region, China (CUHK 24615523), and National Natural Science Foundation of China (No. 12301613). Zijia Wang acknowledges the support from internal grants from The Chinese University of Hong Kong and National Natural Science Foundation of China (No. 12401623).
Appendix A Technical Details
A.1 Plateau Death Models
@@
The remaining steps follow exactly the same as before, and we will end up with the same survival function as in equation (3).
@@
If $a=1$, $p=\alpha$, and $b=\frac{v}{\phi}$, then we recover the same survival function as in equation (3).
@@ A.2 Alternative Expression of Mortality Laws
A.3 Vitality Models with Dynamic Parameters